### v1

@@ 3.2 Infinite-Horizon Case
While the DRKF in Theorem˜1 provides a tractable solution to the finite-horizon DRSE problem, our focus is on its long-term behavior. To study this regime, we assume that both the system dynamics and the nominal uncertainty statistics are time invariant.
@@
The system (1) is time invariant, i.e., $A_{t}\equiv A,C_{t}\equiv C$. Moreover, the nominal uncertainty distributions are stationary, so that $\hat{w}_{t}\equiv\hat{w},\hat{v}_{t}\equiv\hat{v}$ and $\hat{\Sigma}_{v,t}\equiv\hat{\Sigma}_{v},\hat{\Sigma}_{w,t}\equiv\hat{\Sigma}_{w}$ for all $t\geq 0$, where $\hat{\Sigma}_{w}\in\mathbb{S}^{n_{x}}_{++}$ and $\hat{\Sigma}_{v}\in\mathbb{S}^{n_{y}}_{++}$.
Under Assumption˜2, the stage-wise SDPs (5)–(7) reduce to a time-invariant recursion for the covariance matrices. The resulting sequences of least-favorable noise covariances $\{\Sigma_{w,t}^{*},\Sigma_{v,t}^{*}\}$ and state covariances $\{\Sigma_{x,t}^{-},\Sigma_{x,t}\}$ may converge to limiting values $\{\Sigma_{w,\infty},\Sigma_{v,\infty}\}$ and $\{\Sigma_{x,\infty}^{-},\Sigma_{x,\infty}\}$ as $t\to\infty$. Whether such convergence holds depends on additional conditions, which will be established later in Section˜4.2. Motivated by this possibility, we seek a steady-state DR estimator with constant gain.
In the infinite-horizon setting, the DRSE problem reduces to finding these steady-state covariances and the corresponding steady-state DR Kalman gain. Since the dynamics and nominal statistics are fixed, the stage-wise SDP (5) simplifies to the following time-invariant convex program:
@@
As in (7), the optimization problem above can be reformulated as an equivalent steady-state SDP:
@@
The optimal solution $\{\Sigma_{w,\infty}^{*},\Sigma_{v,\infty}^{*},\Sigma_{x,\infty}^{-,*},\Sigma_{x,\infty}^{*}\}$ defines the least-favorable stationary covariances, from which the corresponding steady-state DR Kalman gain is
@@
The resulting time-invariant state estimator is
@@
Compared with the finite-horizon DRKF, the steady-state filter (15) requires solving only a single stationary SDP offline. This makes it well-suited for real-time implementation. More importantly, it optimizes directly for worst-case asymptotic performance, ensuring robustness to persistent disturbances and long-term distributional errors. These properties are particularly valuable in safety-critical systems with slow or unstable dynamics.
3.3 Algorithms
We now summarize the proposed DRKF methods for both the time-varying and steady-state settings. These algorithms follow directly from the tractable reformulations in Section˜3 and the recursive structure established in Theorem˜1. Although both versions share the same online recursion as the classical KF, they differ in how the covariance matrices, and hence the Kalman gains, are computed to ensure robustness against distributional ambiguity.
3.3.1 Time-Varying DRKF
As shown in Algorithm˜1, the time-varying DRKF is implemented by first solving the stage-wise SDP problems (4) and (5) offline for each time step. If the nominal covariance matrices are known a priori, the least-favorable noise covariances and the DR gains $\{K_{t}\}$ can be precomputed offline for a given horizon $T$. At each stage, the gain $K_{t}$ is computed using the least-favorable prior state covariance and the least-favorable measurement noise covariance. The online stage then consists only of the measurement update and state prediction steps (9) and (11). Upon receiving a new measurement $y_{t}$, the update and prediction steps are carried out exactly as in the classical KF, except that they use the DR gains computed offline and the nominal noise means.
@@ 3.3.1 Time-Varying DRKF
2:Solve (4) to obtain $(\Sigma_{x,0}^{-*},\Sigma_{v,0}^{*})$ $\triangleright$ Offline stage
3:Compute DR gain $K_{0}$ via (8)
@@
3.3.2 Steady-State DRKF
Under the time-invariant assumptions of Section˜3.2, the DRKF recursion may converge to a stationary solution. In this setting, the procedure in Algorithm˜1 simplifies considerably. Rather than solving an SDP at every stage, one solves the single stationary SDP (13) offline to obtain the least-favorable noise covariances, the steady-state prior and posterior covariances, and the resulting steady-state DR Kalman gain $K_{\infty}^{*}$. The online recursion is identical in structure to the classical steady-state KF, making the approach particularly suitable for long-horizon or real-time applications. The complete procedure is given in Algorithm˜2.
@@ 3.3.2 Steady-State DRKF
In the steady state regime, ambiguity in the initial state distribution becomes irrelevant: the Wasserstein radius $\theta_{x_{0}}$ affects only the transient phase and does not influence the asymptotic solution. As discussed in Section˜4.2 and Remark˜5, this dependence disappears after finitely many steps. Thus, the steady-state DRKF depends only on $(\theta_{w},\theta_{v})$ and the nominal noise covariances.
3.4 Discussion on Computational Complexity
The main computational burden in both DRKF variants arises in the offline stage, where SDPs must be solved. The online recursion has the same per-step complexity as the classical KF.
@@ 3.4 Discussion on Computational Complexity
For comparison, the stage-wise DRKF of [16] and the DR-MMSE estimator of [17] also require solving a convex program at each stage. The former uses the joint state-measurement covariance of dimension $(n_{x}+n_{y})$, leading to similar polynomial scaling but with larger LMI blocks and constants. Like our time-varying DRKF, these methods scale linearly in $T$; however, neither admits a steady-state variant that reduces the offline computation to a single SDP.
Temporally coupled ambiguity sets, as in [20], can model long-range dependence across the noise sequence but incur substantially higher computational cost: all $T$ stages are merged into a single horizon-level SDP whose number of variables and LMI block sizes both grow linearly with $T$. The resulting interior-point complexity, $\tilde{\mathcal{O}}(n_{x}^{6}T^{6})$, quickly becomes prohibitive even for moderate horizons. By contrast, our noise-centric formulation solves a fixed-size SDP at each stage, so the offline cost grows only linearly in $T$.
In the infinite-horizon setting, [20] characterizes the steady-state filter via a frequency-domain max-min problem over Toeplitz operators, producing a generally non-rational filter that requires spectral factorization and approximation for implementation. No finite-dimensional formulation is available. By contrast, our steady-state DRKF is obtained from a single stationary SDP of size $(n_{x}+n_{y})$ with offline complexity $\tilde{\mathcal{O}}\left((n_{x}+n_{y})^{3.5}\right)$, directly yielding a constant gain $K_{\infty}^{*}$ and an online recursion matching the classical steady-state KF. Although this stage-wise ambiguity does not model temporal noise correlations, it preserves the separability of the Riccati recursion and keeps the DR filter tractable for real-time use.
4 Theoretical Analyses
Having established the DRKF formulations in both the finite- and infinite-horizon settings, we now analyze their fundamental theoretical properties. Our goal is to show that DRKF retains key structural guarantees of the classical KF, such as stability and boundedness, while providing provable robustness under distributional uncertainty. In particular, we demonstrate that the DRKF covariance recursion remains spectrally bounded and that its steady-state solution inherits standard convergence properties of the classical filter.
@@ 4.1 Spectral Boundedness
Our starting point is a well-known characterization of the Bures–Wasserstein distance.
@@
[33, Theorem 1] Let $A,B\in\mathbb{S}^{n}_{+}$. The Bures–Wasserstein distance satisfies
@@
where $O(n)$ denotes the orthogonal group. The minimizer exists and is given by the unitary factor in the polar decomposition of $B^{1/2}A^{1/2}$.
@@
The proof is provided in Appendix B.1. This result shows that Bures–Wasserstein balls impose both lower and upper eigenvalue bounds on any feasible covariance matrix. Equivalently, the Bures–Wasserstein ball defines an eigenvalue tube around the nominal covariance. This geometric interpretation is essential: it guarantees that our DRKF cannot produce arbitrarily inflated or degenerate covariances, even under worst-case perturbations of the noise distributions allowed by the ambiguity sets. In contrast, KL-based ambiguity sets can admit pathological least-favorable distributions [29].
@@
$\begin{split}\underline{\lambda}_{v,t}&:=\lambda_{\min}(\hat{\Sigma}_{v,t}),\qquad\overline{\lambda}_{v,t}:=\overline{\lambda}(\hat{\Sigma}_{v,t},\theta_{v}),\\ \underline{\lambda}_{w,t}&:=\lambda_{\min}(\hat{\Sigma}_{w,t}),\qquad\overline{\lambda}_{w,t}:=\overline{\lambda}(\hat{\Sigma}_{w,t},\theta_{w})\\ \underline{\lambda}_{x,0}&:=\lambda_{\min}(\hat{\Sigma}_{x,0}^{-}),\qquad\overline{\lambda}_{x,0}:=\overline{\lambda}(\hat{\Sigma}_{x,0}^{-},\theta_{x_{0}}).\end{split}$ (17)
Then, for all $t>0$, the least-favorable covariances $\Sigma_{v,t}^{*},\Sigma_{w,t}^{*}$ and $\Sigma_{x,0}^{-,*}$ obtained from (4)–(5) satisfy
@@
The proof is provided in Appendix B.2. In the time-invariant setting, the constants $\underline{\lambda}_{v},\overline{\lambda}_{v},\underline{\lambda}_{w},\overline{\lambda}_{w}$ are fixed. Thus, Corollary˜2 ensures that the DRKF operates within a compact spectral envelope determined solely by the nominal statistics and the Wasserstein radii.
@@
Consider two classical KFs with noise covariances chosen according to the lower and upper bounds in (17). Let $(\underline{\Sigma}_{x,t}^{-},\underline{\Sigma}_{x,t})$ and $(\overline{\Sigma}_{x,t}^{-},\overline{\Sigma}_{x,t})$ denote their respective prior and posterior error covariances, with initial conditions $\underline{\Sigma}_{x,0}^{-}=\underline{\lambda}_{x,0}I$ and $\overline{\Sigma}_{x,0}^{-}=\overline{\lambda}_{x,0}I$. Then, for all $t\geq 0$, the DRKF covariances $(\Sigma_{x,t}^{-},\Sigma_{x,t})$ satisfy
@@ 4.2 Convergence
Having established the key theoretical properties of the least-favorable covariance matrices, we now derive conditions under which the DRKF admits a unique steady-state solution, and the sequence of time-varying gains converges to its steady-state value. Specifically, recall that the steady-state DR Kalman gain $K_{\infty}$ is defined in (14). Our goal is to show that
$\lim_{t\to\infty}K_{t}=K_{\infty}.$
Under the standing assumptions, the steady-state SDP (13) has at least one optimal solution, since its objective is continuous and the feasible set is nonempty and compact due to the Bures-Wasserstein spectral bounds. Similarly, the time-varying SDP problems (4) and (5) admit optimal solutions, and therefore the sequence of DR Kalman gains $\{K_{t}\}$ is well defined.
@@
with $\Sigma_{x,0}^{-}=\Sigma_{x,0}^{-,*}$. Since $\Sigma_{x,0}^{-}\succeq\lambda_{\min}(\hat{\Sigma}_{x,0}^{-})I_{n_{x}}$ and $\Sigma_{x,t}^{-}\succeq\lambda_{\min}(\hat{\Sigma}_{w})I_{n_{x}}$ for $t>0$ at the optimum, $\Sigma_{x,t}^{-}$ is positive definite and the recursion is well defined. This motivates the following definition of the DR Riccati mapping:
@@
As in the classical KF, controllability and observability are needed to guarantee a unique steady state.
@@
The pair $(A,\hat{\Sigma}_{w}^{1/2})$ is controllable and the pair $(A,C)$ is observable.
To isolate the effect of distributional robustness in the process channel, we parameterize the perturbation through $\Delta_{t}$.
Assumption 4.
There exists $\Delta_{t}\in\mathbb{S}^{n_{x}}_{+}$ such that $A\Delta_{t}A^{\top}=\Sigma_{w,t}^{*}-\hat{\Sigma}_{w}$.
Under Assumption˜4, the recursion (18) can be rewritten as
$r_{\mathbb{D}}(\Sigma_{x,t}^{-})=A\bigl((\Sigma_{x,t}^{-})^{-1}+C^{\top}\hat{\Sigma}_{v}^{-1}C-\Phi_{t}\bigr)^{-1}A^{\top}+\hat{\Sigma}_{w},$ (20)

### v3

@@ 3.2 Infinite-Horizon Case
While the DRKF in Theorem˜1 provides a tractable solution to the finite-horizon DRSE problem, our primary interest lies in its long-term behavior. To study this regime, we assume that both the system dynamics and the nominal uncertainty statistics are time-invariant.
@@
The system (1) is time-invariant, i.e., $A_{t}\equiv A,C_{t}\equiv C$. Moreover, the nominal uncertainty distributions are stationary, so that $\hat{w}_{t}\equiv\hat{w},\hat{v}_{t}\equiv\hat{v}$ and $\hat{\Sigma}_{v,t}\equiv\hat{\Sigma}_{v},\hat{\Sigma}_{w,t}\equiv\hat{\Sigma}_{w}$ for all $t\geq 0$, where $\hat{\Sigma}_{w}\in\mathbb{S}^{n_{x}}_{++}$ and $\hat{\Sigma}_{v}\in\mathbb{S}^{n_{y}}_{++}$. The nominal initial state covariance satisfies $\hat{\Sigma}_{x,0}^{-}\in\mathbb{S}^{n_{x}}_{++}$.
Under Assumption˜2, the stage-wise SDPs (5)–(7) reduce to a time-invariant recursion for the covariance matrices. The resulting sequences of least-favorable noise covariances $\{\Sigma_{w,t}^{*},\Sigma_{v,t}^{*}\}$ and state covariances $\{\Sigma_{x,t}^{-},\Sigma_{x,t}\}$ may converge to limiting values $\{\Sigma_{w,\infty},\Sigma_{v,\infty}\}$ and $\{\Sigma_{x,\infty}^{-},\Sigma_{x,\infty}\}$ as $t\to\infty$. Whether such convergence holds depends on additional conditions, which are established later in Section˜4.2. Motivated by this, we seek a steady-state DR estimator with a constant gain.
In the infinite-horizon setting, the DRSE problem reduces to computing these steady-state covariances and the associated DR Kalman gain. Since the dynamics and nominal statistics are time-invariant, the stage-wise SDP (5) simplifies to the following convex program:
@@
As in the finite-horizon case, this problem admits an equivalent SDP formulation. In particular, it can be written as the following single stationary SDP:
@@
The optimal solution $\{\Sigma_{w,\infty}^{*},\Sigma_{v,\infty}^{*},\Sigma_{x,\infty}^{-,*},\Sigma_{x,\infty}^{*}\}$ defines the least-favorable stationary covariances and yields the steady-state DR Kalman gain
@@
The resulting time-invariant estimator is
@@
Compared with the finite-horizon DRKF, the steady-state filter (15) requires solving only a single stationary SDP offline. This makes it particularly suitable for real-time implementation. More importantly, it directly optimizes worst-case asymptotic performance, providing robustness to persistent disturbances and long-term distributional errors–an essential property for safety-critical systems with slow or unstable dynamics.
@@
2:Solve (4) to obtain $(\Sigma_{x,0}^{-,*},\Sigma_{v,0}^{*})\triangleright$ Offline stage
3:Compute DR gain $K_{0}$ via (8) and update the posterior covariance $\Sigma_{x,0}$ via (10)
@@
3.3 Algorithms and Complexity
We summarize the proposed DRKF methods for both the time-varying and steady-state settings. The two variants differ only in how the noise covariances (and hence the Kalman gains) are computed offline to ensure robustness against distributional ambiguity.
3.3.1 Time-Varying and Steady-State DRKF
As shown in Algorithm˜1, the time-varying DRKF is implemented by solving the stage-wise SDP problems (4) and (5) offline over a horizon $T$. This yields the least-favorable noise covariances and the corresponding DR gains $\{K_{t}\}$. If the nominal covariances are known a priori, these gains can be fully precomputed.
Under the time-invariant assumptions of Section˜3.2, the DRKF recursion may converge to a stationary solution. In this case, the offline stage reduces to solving the single stationary SDP (13), which yields the least-favorable noise covariances, the steady-state prior and posterior covariances, and the constant DR Kalman gain $K_{\infty}^{*}$.
In both cases, the online stage performs only the standard KF measurement update and prediction, and thus has the same per-step computational complexity as the classical KF.
3.3.2 Discussion on Computational Complexity
The computational cost of the proposed DRKF methods is dominated by the offline SDP computations; the online recursion has the same per-step complexity as the classical KF.
@@ 3.3.2 Discussion on Computational Complexity
For comparison, the stage-wise DRKF of [26] and the DR-MMSE estimator of [23] also require solving a convex program at each stage. The former relies on joint state–measurement covariances of dimension $(n_{x}+n_{y})$, leading to similar polynomial scaling but with larger LMI blocks and constants. Like our time-varying DRKF, these methods scale linearly with $T$, but they do not admit a steady-state formulation that reduces the offline computation to a single SDP.
Temporally coupled ambiguity sets, as in [13], model long-range dependence across the noise sequence at the cost of significantly higher complexity: all $T$ stages are merged into a single horizon-level SDP whose size grows linearly with $T$, resulting in interior-point complexity of $\tilde{\mathcal{O}}(n_{x}^{6}T^{6})$. By contrast, our noise-centric formulation solves a fixed-size SDP at each stage, so the offline cost grows only linearly in $T$.
In the infinite-horizon setting, [13] characterizes the steady-state filter via a frequency-domain max-min problem over Toeplitz operators, producing a generally non-rational filter that requires spectral factorization and approximation for implementation. No finite-dimensional formulation is available. In contrast, our steady-state DRKF is obtained from a single stationary SDP of size $(n_{x}+n_{y})$ with offline complexity $\tilde{\mathcal{O}}\left((n_{x}+n_{y})^{3.5}\right)$. It directly yields a constant gain $K_{\infty}^{*}$ and an online recursion matching the classical steady-state KF. Although this stage-wise ambiguity does not model temporal noise correlations, it preserves the separability of the Riccati recursion and keeps the DR filter tractable for real-time use.
4 Theoretical Analysis
Having established the DRKF formulations in both the finite- and infinite-horizon settings, we now analyze their fundamental theoretical properties. Our goal is to show that the DRKF retains key structural guarantees of the classical KF, such as stability and boundedness, while providing provable robustness under distributional uncertainty. In particular, we demonstrate that the DRKF covariance recursion remains spectrally bounded and that its steady-state solution inherits standard convergence properties of the classical filter.
@@ 4.1 Spectral Boundedness
Our starting point is a well-known characterization of the (non-squared) Bures–Wasserstein distance.
@@
[2, Thm. 1] Let $A,B\in\mathbb{S}^{n}_{+}$. The Bures–Wasserstein distance satisfies
@@
where $O(n)$ denotes the orthogonal group. The minimizer exists and is given by the orthogonal factor in the polar decomposition of $B^{1/2}A^{1/2}$.
@@
The proof is provided in Appendix B.1. This result shows that a Bures–Wasserstein ball induces both upper and lower spectral bounds on feasible covariance matrices, defining an eigenvalue tube around the nominal covariance. This geometric perspective is essential: it rules out pathological behavior in which worst-case covariances can become arbitrarily inflated or collapse in directions that are inconsistent with the nominal statistics. As a result, the DRKF operates within a controlled spectral envelope under distributional perturbations allowed by the ambiguity sets. By contrast, KL-based ambiguity sets may admit least-favorable distributions with highly distorted or ill-conditioned covariance structure, leading to potentially pathological estimators [5].
@@
$\begin{split}\underline{\lambda}_{v,t}&:=\lambda_{\min}(\hat{\Sigma}_{v,t}),\qquad\overline{\lambda}_{v,t}:=\overline{\lambda}(\hat{\Sigma}_{v,t},\theta_{v}),\\ \underline{\lambda}_{w,t}&:=\lambda_{\min}(\hat{\Sigma}_{w,t}),\qquad\overline{\lambda}_{w,t}:=\overline{\lambda}(\hat{\Sigma}_{w,t},\theta_{w}),\\ \underline{\lambda}_{x,0}&:=\lambda_{\min}(\hat{\Sigma}_{x,0}^{-}),\qquad\overline{\lambda}_{x,0}:=\overline{\lambda}(\hat{\Sigma}_{x,0}^{-},\theta_{x_{0}}).\end{split}$ (17)
Then, the least-favorable covariances $\Sigma_{v,t}^{*},\Sigma_{w,t}^{*}$ and $\Sigma_{x,0}^{-,*}$ obtained from (4)–(5) satisfy
@@
The proof is provided in Appendix B.2. In the time-invariant setting, the constants $\underline{\lambda}_{w}\coloneq\lambda_{\min}(\hat{\Sigma}_{w})$, $\underline{\lambda}_{v}\coloneq\lambda_{\min}(\hat{\Sigma}_{v})$, $\overline{\lambda}_{w}\coloneq\overline{\lambda}(\hat{\Sigma}_{w},\theta_{w})$, and $\overline{\lambda}_{v}\coloneq\overline{\lambda}(\hat{\Sigma}_{v},\theta_{v})$ are fixed. Thus, Corollary˜2 ensures that the DRKF operates within a compact spectral envelope determined solely by the nominal statistics and the Wasserstein radii.
@@
Consider two classical KFs with noise covariances chosen according to the lower and upper bounds in (17). Let $(\underline{\Sigma}_{x,t}^{-},\underline{\Sigma}_{x,t})$ and $(\overline{\Sigma}_{x,t}^{-},\overline{\Sigma}_{x,t})$ denote their respective prior and posterior error covariances, with initial conditions $\underline{\Sigma}_{x,0}^{-}=\underline{\lambda}_{x,0}I_{n_{x}}$ and $\overline{\Sigma}_{x,0}^{-}=\overline{\lambda}_{x,0}I_{n_{x}}$. Then, for all $t\geq 0$, the DRKF covariances $(\Sigma_{x,t}^{-},\Sigma_{x,t})$ satisfy
@@ 4.2 Convergence
Having established the key theoretical properties of the least-favorable covariance matrices, we now derive conditions under which the DRKF admits a unique steady-state solution, and the sequence of time-varying gains converges to its steady-state value. Specifically, recall that the stationary DR Kalman gain $K_{\infty}^{*}$ is defined in (14). Our goal is to show that
$\lim_{t\to\infty}K_{t}=K_{\infty}^{*}.$
Under the standing assumptions, the time-varying SDP problems (4) and (5) admit optimal solutions (the feasible sets are nonempty and compact under the Bures–Wasserstein bounds), and therefore the sequence of DR Kalman gains $\{K_{t}\}$ is well defined. In the time-invariant setting, the steady-state SDP (13) is feasible (e.g., by choosing the nominal covariances), and below we show that the DR Riccati recursion converges to a unique fixed point characterized by this SDP.
@@
with $\Sigma_{x,0}^{-}=\Sigma_{x,0}^{-,*}$. If $\hat{\Sigma}_{x,0}^{-}\succ 0$, then $\Sigma_{x,0}^{-}\succ 0$ and the recursion is well defined. [^8] If $\hat{\Sigma}_{x,0}^{-}\succeq 0$, the uniform bound $\Sigma_{w,t}^{*}\succeq\underline{\lambda}_{w}I$ ensures that $\Sigma_{x,1}^{-}\in\mathbb{S}^{n_{x}}_{++}$, so the information-form recursion is well defined from time $t=1$ onward. This motivates the following definition of the DR Riccati mapping:
@@
Under Assumption˜2, the problem data of the stage-wise SDP are time-invariant. Accordingly, the sequence $\{(\Sigma_{w,t}^{*},\Sigma_{v,t}^{*})\}$ is generated by repeatedly applying the same SDP template along the state covariance trajectory. Observing the structural similarity with Riccati equations arising in robust and risk-sensitive filtering (e.g., [32, 10]), we pursue contraction-based analysis in this subsection.
We begin by imposing the standard observability assumption. [^9] Since $\hat{\Sigma}_{w}\succ 0$ and $\Sigma_{w,t}^{*}\succeq\underline{\lambda}_{w}I$, the process noise covariance is uniformly positive definite; hence no additional controllability assumption is required.
@@
The pair $(A,C)$ is observable.
To establish strict contraction under the observability assumption, we employ a downsampled (lifted) representation of the DR Riccati recursion, since the one-step mapping may fail to be contractive due to rank-deficient measurement operators, whereas observability guarantees full column rank of the lifted observability matrix.
Fix $N\in\mathbb{N}$ and consider the system (1) under the least-favorable distributions. Define the downsampled state as $x_{t}^{d}\coloneq x_{tN}$. Downsampling yields an NN-block lifted model with block-diagonal uncertainty structure, enabling a contraction argument similar to [34, 36]. Let $\bm{w}_{t}^{N}\coloneqq[w_{tN+N-1}^{\top},\dots,w_{tN}^{\top}]^{\top}$, and $\bm{v}_{t}^{N}\coloneqq[v_{tN+N-1}^{\top},\dots,v_{tN}^{\top}]^{\top}$ denote the stacked process and measurement noise vectors, with means $\mathbf{1}_{N}\otimes\hat{w}$ and $\mathbf{1}_{N}\otimes\hat{v}$, and covariances $\Sigma_{w,t}^{N,*}\coloneq\operatorname{blkdiag}(\Sigma_{w,tN+N-1}^{*},\dots,\Sigma_{w,tN}^{*})$ and $\Sigma_{v,t}^{N,*}\coloneq\operatorname{blkdiag}(\Sigma_{v,tN+N-1}^{*},\dots,\Sigma_{v,tN}^{*})$, respectively. The downsampled dynamics then take the form
$\begin{split}x_{t+1}^{d}&=A^{N}x_{t}^{d}+\mathcal{R}_{N}\bm{w}_{t}^{N}\\ \bm{y}_{t}^{N}&=\mathcal{O}_{N}x_{t}^{d}+\mathcal{H}_{N}\bm{w}_{t}^{N}+\bm{v}_{t}^{N},\end{split}$
where $\bm{y}_{t}^{N}\coloneqq[y_{tN+N-1}^{\top},\dots,y_{tN}^{\top}]^{\top}$ is the stacked observation vector, while the block reachability and observability matrices are defined as
$\mathcal{R}_{N}$ $\coloneqq\left[I_{n_{x}},\,A,\,\ldots,\,A^{N-1}\right]$ (20)
$\mathcal{O}_{N}$ $\coloneqq\left[(CA^{N-1})^{\top},\,\ldots,\,(CA)^{\top},\,C^{\top}\right]^{\top}.$ (21)
The block Toeplitz matrix $\mathcal{H}_{N}\in\real{Nn_{y}\times Nn_{x}}$ captures the propagation of process noise into future outputs and is defined elementwise by $\mathcal{H}_{N}]_{ij}=CA^{j-i-1},j>i$, and zero otherwise, for $i,j=1,\ldots,N$. Let $\bm{u}_{t}^{N}\coloneq\mathcal{H}_{N}\bm{w}_{t}^{N}+\bm{v}_{t}^{N}$. Its covariance is given by $\Sigma_{u,t}^{N}=\mathcal{H}_{N}\Sigma_{w,t}^{N,*}\mathcal{H}_{N}^{\top}+\Sigma_{v,t}^{N,*}$. Adopting the procedure in [16] and denoting the downsampled prior covariance matrix by $\Sigma_{x,t}^{-,d}:=\Sigma_{x,tN}^{-}$, the DR Riccati recursion associated with the downsampled model becomes
$\Sigma_{x,t+1}^{-,d}=r_{\mathbb{D},t}^{d}(\Sigma_{x,t}^{-,d}),$ (22)